money is borrowed at 12% simple interest. after 1 year, $896 pays off the loan.
how much was borrowed originally?

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

896/1.12=$800

Answer by stanbon(75887)

money is borrowed at 12% simple interest. after 1 year, $896 pays off the loan.
how much was borrowed originally?
-----------------------------
A(t) = P(1+rt)
---
896 = P(1+0.12*1)
---
896 = P(1.12)
---
Principal = 896/1.12
Principal = $800
======================
Cheers,
Stan H.
======================